OpenStudy (anonymous):

find the principal (P) if I=$1260,R=4.5%, T=7 years please help

OpenStudy (jdoe0001):

i = prt you want to find "p" or Principal, just solve for "P" then plug in the values given

OpenStudy (jdoe0001):

well... .keep in mind that 4.5% is has to be entered in decimals one sec

OpenStudy (jdoe0001):

\(\bf i={\color{brown}{ p}}rt\\ i=interest\to 1,260\\ r=rate\to 4.5\%\to \frac{4.5}{l00}\to 0.045\\ t=years\to 7 \\ \quad \\ i={\color{brown}{ p}}rt\implies 1,260={\color{brown}{ p}}\cdot 0.045t\) solve for "p" to find the Principal amount, or original deposit

OpenStudy (jdoe0001):

hmm kinda missed the years there... so anyhow \(\bf i={\color{brown}{ p}}rt\\ i=interest\to 1,260\\ r=rate\to 4.5\%\to \frac{4.5}{l00}\to 0.045\\ t=years\to 7 \\ \quad \\ i={\color{brown}{ p}}rt\implies 1,260={\color{brown}{ p}}\cdot 0.045\cdot 7\)